OpenStudy (anonymous):

Find a quadratic equation with roots –1 + 4i and –1 – 4i. Show all your work

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

(x- ( -1+4i ) )*(x-( -1-4i) )

OpenStudy (anonymous):

(x- ( -1+4i ) )(x-( -1-4i) ) = x^2-(-1-4i)-(-1+4i)+(-1+4i)(-1-4i)= ?

OpenStudy (anonymous):

okay I think I figured it out, is the answer x^2 + 2x + 17 = 0

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

okay thank you both